# Learning in Deep Factor Graphs with Gaussian Belief Propagation

### v1

@@ 3.2 Continual Learning
We now describe how we can do continual learning of our model parameters with Bayesian filtering. For generality, we use $\mathbf{\Psi}=\{\mathbf{\psi}_{l}\}_{l=1}^{L}$ to denote the set of all parameters where $\mathbf{\psi}_{l}$ is the vector of parameters for layer $l$. After initialising parameter priors $p_{t=1}\left(\mathbf{\psi}_{l,i}\right)\leftarrow\mathcal{N}\left(0,\sigma\right)$ we perform the following for each task $t$ in the sequence of datasets $[\mathbf{z}_{1},\ldots,\mathbf{z}_{T}]$:
@@
    construct a new graph with task dataset $\mathbf{z}_{t}$,
@@
3.3 Efficient Gaussian Belief Propagation
3.3.1 Factor to Variable Message Optimisation
Efficient inference is a prerequisite to our models being applicable in practice. However, the inversion of a $\left(V_{j}-1\right)\times\left(V_{j}-1\right)$ matrix to compute $\Sigma_{\setminus i,\setminus i}^{\left(\phi_{j}+m\right)}$ in (8) is $O\left((V_{j}-1)^{3}\right)$ complexity, bottlenecking GBP. To alleviate this, we exploit the structure of matrix being inverted. In particular, i) for factors with observation dimension $M=\dim\left(\mathbf{y}\right)<V_{j}$, the precision $\Lambda^{(\phi_{j})}$ (9) is low-rank, and ii) for graphs with scalar variable nodes, $\mathbf{D}$ is diagonal. Thus their sum may be efficiently inverted via the Woodbury identity (woodbury1950inverting). Further savings come from reusing intermediates when computing messages to multiple variables (9). These optimisations change the complexity of updating messages from a factor to all $V$ variables, from $O(V\left(V-1\right)^{3})$ to $O\left(VM^{3}\right)$. Space complexity is changed from $O(V^{2})$ to $O(VM+M^{2})$. Both are significant savings when $M<<V$. See App. A for details.
3.3.2 Implementation
Our factor graphs have regular, repeated structure making our GBP amenable to GPU parallelisation. We implement our method in TensorFlow (Tensorflow:MISC2015) to make use of GPU support and efficient linear algebra operations. For a given batch, we initialise the graph and update the messages by sweeping forward and backward through the layers: first the layer nearest the input observations, then progressively deeper to the deepest layer and back again. We repeat these sweeps for a specified number of iterations. When updating the messages within each layer, we parallelise all factor to variable messages, and likewise for the variable to factor updates of each variable type (inputs, outputs, parameters as applicable). This message schedule is likely suboptimal, but we leave further exploration for future work. We find that applying damping (murphy2013loopy) and dropout to the factor to variable message updates is sufficient for stable GBP.
For supervised learning experiments, the output observation factors are only included at training time. At test time, we predict by running GBP in the test graph and taking the last layer variables as regression predictions or logits. As our factor graph is a joint model of both inputs and labels we could, in theory, continue to learn from the input distribution at test time. In practice we find performance is slightly better when we fix the learnt parameters, and run GBP on the input, output and activations.
@@ 4 Results
We demonstrate the benefits of our approach with three experiments: some small regression tasks, sequential video denoising and continual learning for image classification. The corresponding code is made available111anonymous.4open.science/r/gbp-learning-anon-F2CB/.
@@
Figure 2: GBP learning in MLP-like factor graphs (1(a)) can solve nonlinear regression and classification tasks.
@@ 4.1 Toy Experiments
We start by verifying that our model can solve tasks requiring nonlinear modelling. To this end, we run GBP learning with single hidden layer, MLP-like factor graphs (Fig. 1(a)) in two settings. The first is “Exclusive-OR”, which is a well-known minimal test for nonlinear modelling (Fig. 1(b)). The second is a nonlinear regression problem (Fig. 1(c)) with $90$ training points. In both cases, we employ a Leaky ReLU activation in the first dense layer. Full details are provided in App. B. It is clear that GBP learning can solve both tasks, confirming that the linearisation method described in Section 2.3 is sufficient to capture nonlinear dependencies.
@@ 4.2 Video Denoising
Figure 3: Video denoising results. Factor graphs with learnable components outperform a hand-specified pairwise smoother with robust factors, and continual learning of the parameters throughout the video further improves performance over per-frame learning. Shaded regions cover ±$\pm 2$ standard errors (SEs) over $10$ seeds.
We now ask whether the learnable components in our model can improve performance over a hand-designed solver. We apply our method to the task of denoising the “bear” video from the DAVIS dataset [^2] Creative Commons Attributions 4.0 License (perazzi2016benchmark), downsampled to $258\times 454$ with bilinear interpolation. We randomly select $10\%$ of the pixels in each frame of the video, and replace their intensities with noise independently sampled from $U(0,1)$. Performance is assessed by how well the denoised image matches the ground truth according to the peak signal-to-noise ratio (PSNR).
We run GBP learning with two models: i) a single transposed convolution layer with four filters, and ii) a five layer model comprising stacked transposed convolution and upsampling layers (see App. C). Both the pixel observation and reconstruction factors have robust energies (see Section 5 of Davison:Ortiz:ARXIV2019 for details) to enable the pixel variables to “switch” between being explained by the noisy observation or by the reconstruction model. As a baseline, we compare to a classical pairwise smoothing model in which neighbouring pixels are encouraged to have similar intensities through smoothness factors shared between adjacent pixels (see e.g ortiz2021visual). Robust energies for these factors aid the preservation of edges present in the original image. The baseline also includes robust energies for pixel observation factors, as in our model.
In both pairwise smoother and GBP learning, we do inference with GBP. Note that in our models, this is jointly estimating the true pixel intensities and the model parameters. We emphasise that our model learns image structure from the noisy images only and without supervision. It is incentivised to do so by the reconstruction factors in noiseless regions. We evaluate two routines for training the parameters: i) learn them from scratch on each frame and ii) do filtering on the parameters over the sequence of frames, with the approach described in Section 3.2. The hyperparameters for all models were tuned using the first five frames of the video. Further details of the final models can be found in App. C. Denoising the entire $82$-frame video with the single layer model took $\sim 8$mins on a NVIDIA RTX 3090 GPU, and the five layer model took $\sim 27$mins. Pairwise smoothing took $\sim 2$mins.
@@
There is clear benefit of depth, with significantly higher PSNR scores for the five layer model (orange/green dotted) compared to the single layer model (orange/green dash). Despite the single layer model having only four filters, we find that it overfits during continual learning. To counteract this, we set the prior for each frame to be an interpolation between the previous posterior and the original prior (plotted; see App. C). In contrast, the deeper model does not require additional regularisation when learning continually. These results imply i) deep factor graphs can have better inductive biases, ii) GBP learning is effective in finding aligned multi-layer representations with only local message updates.
@@ 4.3 Image Classification
Next, we assess our method in a supervised learning context, evaluating it on MNIST333yann.lecun.com/exdb/mnist/. MNIST has Creative Commons Attribution-Share Alike 3.0 license.. We train a convolutional factor graph with architecture similar to Fig. 1: a convolutional layer (with energy as per (11)), followed by a max-pool (14), dense layer (15) and (at train time) a class observation factor (16) on the output logit variables. Training is minibatched via the Bayesian filtering approach described in Section 3.2, so the model only sees each datapoint once after which it can be discarded. Full details of the model are in App. E.1. To gauge sample efficiency, we train on subsets of varying sizes, as well as the full training set. We compare against a linear classifier baseline which predicts logits via a dense projection of the image. This baseline is trained with gradient descent. Further details of the baseline are included in App. E.2. Both models were tuned on validation sets generated by randomly subsampling $15\%$ of the training set.
We find the GBP Learning in the deep model achieves a test accuracy of $96.3\%$ when trained on the full training set (Fig. 4), significantly higher than the linear classifier ($92.6\%$). This implies that our approach can learn useful features and capture nonlinear relationships. In addition, our method shows better sample efficiency, beating the baseline for every training set size, despite only seeing each datapoint once. In contrast, the linear classifier is trained for multiple epochs. We also trained our model in a layer-wise asynchronous manner by uniformly sampling a layer whose messages to update at each step. Such an approach yields a final accuracy of $96.35\pm 0.03\%$ (±$\pm 1$SE over $3$ seeds), indistinguishable from $96.4\pm 0.2\%$ achieved by the same model trained with synchronous forward/backward sweeps (details in App. E.3). This result implies that GBP Learning can work well when the model is distributed over multiple processors, without the need for global synchronisation. We find the GBP Learning in the deep model achieves a test accuracy of $96.3\%$ when trained on the full training set (Fig. 4), significantly higher than the linear classifier ($92.6\%$). This implies that our approach can learn useful features and capture nonlinear relationships. In addition, our method shows better sample efficiency, beating the baseline for every training set size, despite only seeing each datapoint once. In contrast, the linear classifier is trained for multiple epochs. We also trained our model in a layer-wise asynchronous manner by uniformly sampling a layer whose messages to update at each step. Such an approach yields a final accuracy of $96.35\pm 0.03\%$ (±$\pm 1$SE over $3$ seeds), indistinguishable from $96.4\pm 0.2\%$ achieved by the same model trained with synchronous forward/backward sweeps (details in App. E.3).
@@
Full dataset training with GBP Learning takes around $3$ hours on NVIDIA RTX 3090 GPU. While this is considerably slower than the few minutes it took to train the linear classifier on CPU (and would take to train a small NN), we note that the software to train such models has benefited from decades of optimisation, as have well-established hardware platforms like GPUs. In contrast, our current GBP Learning implementation runs on a hardware and software stack optimised for a different set of models. We believe orders of magnitude efficiency gains could be made by running GBP Learning on a more tailored setup including modern processors with significant on-chip memory. See Section 6 for further discussion.
Figure 4: MNIST results. A deep factor graph trained with GBP learning outperforms a linear model trained with gradient descent, over a pareto front of training set sizes. Error bars cover ±$\pm 2$ SEs over $5$ random seeds.
5 Related Work
Our models can be viewed as probabilistic energy-based models (EBMs lecun2006tutorial; du2019implicit) whose energy functions are the sum of the quadratic energies for all factors in the graph. The benefit of using a Gaussian factor graph is a model which normalises in closed form without resorting to expensive and tricky-to-tune MCMC sampling. While the quadratic energy may seem constraining, we add capacity with the introduction of non-linear factors and overparameterisation. For most EBMs, parameters are viewed as attributes of the factors, where we include them as variables in our graph. As a result, learning and inference are a single combined procedure in our approach, but two distinct stages for other EBMs.
Of the EBM family, restricted Boltzmann machines (RBMs smolensky1986information) are of particular relevance. Their factor graph resembles a single layer, fully connected version of our model, however RBMs are models over discrete variables which are unable capture the statistics of continuous natural images. While exponential family generalisations exist, such as Gaussian-Bernoulli RBMs (welling2004exponential), scaling RBMs to multiple layers remains a challenging problem. To our knowledge there are no working examples of training stacked RBMs jointly, instead they are trained greedily with each layer learning to reconstruct the fixed activations of the previous layer after it has converged (hinton2006fast; hinton2006reducing). As an artefact of this, later layers use capacity in learning the artificial correlations introduced by the previous layers rather than residual correlations present in the data. Global alignment is then found by fine tuning with either backprop (hinton2006reducing) or wake-sleep (hinton2006fast). We have found BP training to work in multiple layer factor graphs without issue.
Our approach also relates to Bayesian DL (neal2012bayesian) in that we seek to infer a distribution over parameters of a deep network. However, common methods to train Bayesian NNs (BNNs) are based on e.g. optimising a variational objective (blundell2015weight; gal2016dropout), the Laplace approximation (mackay1992practical; ritter2018scalable), Hamiltonian Monte Carlo (neal2012bayesian), Langevin-MCMC (zhang2019cyclical). Each of these methods comes with their own benefits and drawbacks but we note two features of our method which distinguish it from existing work in this field. The first is a unified procedure of training and prediction, which are two separate stages in existing methods. The second is our activations are random variables, allowing us to model (and resolve) disagreement between bottom-up and top-down signals.
Much work has focused on augmenting BP with DL (nachmani2016learning; satorras2021neural; yoon2019inference; lazaro2021query). In contrast, our method does DL within a BP framework. george2017generative also propose a hybrid vision system in which visual features and graph structure are learnt in a separate process to the discrete BP routine used to parse scenes (given features and graph). lazaro2016hierarchical is more similar to ours in that parameters are treated as variables in the graphical model and updated with BP. However, their factor graph comprises only binary variables, making it ill-suited to natural images. In addition, max-product BP is used which only finds a MAP solution, and not a full posterior estimate. This precludes the ability to learn incrementally via Bayesian filtering.
6 Conclusions
@@
We have introduced GBP Learning, a method for learning in Gaussian factor graphs. Parameters are included as variables in the graph and learnt using the same BP inference procedure used to estimate all other latent variables. Inter-layer factors encourage representations to be locally consistent, and multiple layers can be stacked in order to learn richer abstractions. Experimentally, we have seen a shallow, learnable model can outperform a hand-crafted method on a video denoising task, with further performance gains coming from stacking multiple layers. We have also applied GBP Learning to train a convolutional factor graph on MNIST. The results show it outperforms a linear classifier both in terms of sample efficiency and peak accuracy, despite only seeing each datapoint once. Both sets of experiments demonstrate the efficacy of our continual learning approach.
@@
While we find these initial results encouraging, we highlight scaling up GBP Learning to bigger, more complex models and datasets as an exciting future direction. Our current implementation is built on software and run on hardware heavily optimised for deep learning. Scaling up GBP Learning requires engineering a bespoke hardware/software system, which can better leverage the distributed nature of GBP inference. In particular, we believe processors with memory local to the compute cores (Graphcore; cerebras) to be a promising platform for efficient BP. On such processors, different sections of the factor graphs could be mapped to different cores. Local message updates, between factors and variables on the same core, would be cheap and could be updated at high frequency, with occasional inter-core communication to ensure alignment between different parts of the model. Low-level GBP primitives, written in e.g. Poplar444graphcore.ai/products/poplar, could be used to ensure the message updates make efficient use of high-bandwidth, local memory. A similar system was applied to solving bundle adjustment problems with GBP, and was found to be around 35×35$\times$35 × faster on IPU (Graphcore) than a commonly used CPU-based solver.
xmain_preprint.bbl
Supplementary Material for Learning in Deep Factor Graphs with Gaussian Belief Propagation
Appendix A Factor to Variable Message Update Optimisation

### v2

@@ 3.4 Continual Learning and Minibatching
We now describe how we can do continual learning of model parameters with Bayesian filtering. For generality, we use $\mathbf{\Psi}=\{\mathbf{\psi}_{l}\}_{l=1}^{L}$ to denote the set of all parameters where $\mathbf{\psi}_{l}$ is the vector of those for layer $l$. After initialising parameter priors $p_{t=1}\left(\mathbf{\psi}_{l,i}\right)\leftarrow\mathcal{N}\left(0,\sigma\right)$ we perform the following for each task $t$ in a sequence of datasets $[\mathbf{z}_{1},\ldots,\mathbf{z}_{T}]$:
@@
    construct a copy of the graph with task dataset $\mathbf{z}_{t}$,
@@
4 Related Work
Our models can be viewed as probabilistic energy-based models (EBMs lecun2006tutorial; du2019implicit) whose energy functions are the sum of the quadratic energies for all factors in the graph. The benefit of using a Gaussian factor graph is a model which normalises in closed form without resorting to expensive MCMC sampling. While the quadratic energy may seem constraining, we add capacity with the introduction of non-linear factors and overparameterisation. For most EBMs, parameters are attributes of the factors, where we include them as variables in our graph. As a result, learning and prediction are the same procedure in our approach, but two distinct stages for other EBMs.
Of the EBM family, restricted Boltzmann machines (RBMs smolensky1986information) are of particular relevance. Their factor graph resembles a single layer, fully connected version of our model, however RBMs are models over discrete variables which are unable capture the statistics of continuous natural images. While exponential family generalisations exist, such as Gaussian-Bernoulli RBMs (welling2004exponential), scaling RBMs to multiple layers remains a challenging problem. To our knowledge there are no working examples of training stacked RBMs jointly, instead they are trained greedily, with each layer learning to reconstruct the activations of the trained and fixed previous layer (hinton2006fast; hinton2006reducing). As an artefact of this, later layers use capacity to learn artificial correlations induced by the early layers, rather than correlations present in the data. Global alignment is then found by fine tuning with either backprop (hinton2006reducing) or wake-sleep (hinton2006fast). We have found GBP Learning to work in multiple layer models without issue.
Our approach also relates to Bayesian DL (neal2012bayesian) as we seek to infer a distribution over parameters of a deep network. Common methods to train Bayesian NNs are based on e.g. variational inference (blundell2015weight; gal2016dropout), the Laplace approximation (mackay1992practical; ritter2018scalable), Hamiltonian Monte Carlo (neal2012bayesian), Langevin-MCMC (zhang2019cyclical). Each of these comes with benefits and drawbacks, but we note two distinguishing features of our method. Computationally, prior Bayesian DL methods rely on backprop, and so inherit its restrictions to distributed and asynchronous training, which do not limit GBP Learning. Further, our activations are random variables, allowing us to model (and resolve) disagreement between bottom-up and top-down signals.
Much work has focused on augmenting BP with DL (nachmani2016learning; satorras2021neural; yoon2019inference; lazaro2021query). In contrast, our method does DL within a BP framework. george2017generative propose a hybrid vision system in which visual features and graph structure are learnt in a separate process to the discrete BP routine used to parse scenes (given features and graph). lazaro2016hierarchical is more similar to ours in that parameters are treated as variables in the graphical model and updated with BP. However, their factor graph comprises only binary variables, making it ill-suited to natural images. In addition, max-product BP is used to find a MAP solution, and not a marginal posterior estimate. This precludes the ability to learn incrementally via Bayesian filtering.
Most relevant to our work is that of lucibello2022deep(lucibello2022deep), who use GBP to train MLPs. Their method relies on analytically derived message updates which only apply to dense architectures, and they focus on models with binary weights and sign activations. In contrast, our approach is straightforward to apply to any network structure without rederivation of messages, assuming the appropriate factor energies can be specified. Further, we focus on continuous weights in direct analogy to NN parameters. Though they similarly minibatch by filtering over parameters, they visit each datapoint multiple times and run a small number of GBP iterations during each visit. After so few iterations, the resulting messages are unlikely to constitute an accurate posterior, and ad-hoc “forgetting” factors are necessary to avoid overcounting data seen multiple times. In contrast, we train with only a single epoch, running GBP to convergence on each batch, enabling straightforward filtering.
@@ 5 Results
We demonstrate our approach with three experiments: some small regression tasks, sequential video denoising and image classification. Our TensorFlow (Tensorflow:MISC2015) implementation is made available222anonymous.4open.science/r/gbp-learning-anon-F2CB/.
@@
Figure 2: GBP Learning in MLP-like factor graphs (1(a)) can solve nonlinear regression and classification tasks. (1(b)) was generated with $8$ hidden units, (1(c)) with $16$ hidden units.
@@ 5.1 Toy Experiments
We start by verifying that our model can solve tasks requiring nonlinear modelling. To this end, we run GBP Learning with single hidden layer, MLP-like factor graphs (Fig. 1(a)) in two settings. The first is “Exclusive-OR”, which is a well-known minimal test for nonlinear modelling (Fig. 1(b)). The second is a nonlinear regression problem (Fig. 1(c)) with $90$ training points. In both cases, we employ a nonlinear activation function in the first dense layer. Full details are provided in App. C. It is clear that GBP Learning can solve both tasks, confirming that the linearisation method described in Section 2.3 is sufficient to capture nonlinear dependencies.
@@ 5.2 Video Denoising
We now ask whether the learnable components in our model can improve performance over a hand-designed solver. We apply our method to the task of denoising the “bear” video from the DAVIS dataset [^3] Creative Commons Attributions 4.0 License (perazzi2016benchmark), downsampled to $258\times 454$ with bilinear interpolation. We sample $10\%$ of the pixels in each frame of the video, replacing their intensities with noise drawn from $U(0,1)$. Performance is assessed by how well the denoised image matches the ground truth under the peak signal-to-noise ratio (PSNR).
To evaluate the impact of model depth, we run GBP Learning with two models: i) a single transposed convolution layer with four filters, and ii) a five layer model comprising stacked transposed convolution and upsampling layers (see App. D). Both the pixel observation and reconstruction factors have robust energies (see Section 5 of Davison:Ortiz:ARXIV2019 for details) to enable the pixel variables to “switch” between being explained by the noisy observation or by the reconstruction model. As a baseline, we compare against a hand-designed method with no learnable parameters. In this pairwise smoothing model, neighbouring pixels are encouraged to have similar intensities via shared smoothness factors (see e.g ortiz2021visual). Robust energies for these factors aid the preservation of edges present in the original image. The baseline also includes robust energies for pixel observation factors, as in our model.
In both the pairwise smoother and GBP Learning, we do inference with GBP. Note that in our models, GBP is jointly estimating the true pixel intensities and the model parameters. We emphasise that our model learns image structure from the noisy images only and without supervision. It is incentivised to do so by the reconstruction factors in noiseless regions. To evaluate our continual learning approach, we use two routines for training the parameters: i) learn them from scratch on each frame and ii) learn them incrementally, by doing filtering on the parameters as described in Section 3.4. The hyperparameters for all models were tuned using the first five frames of the video. Further details of the final models can be found in App. D. Denoising the entire $82$-frame video with the single layer model took $\sim 8$mins on a NVIDIA RTX 3090 GPU, and the five layer model took $\sim 27$mins. Pairwise smoothing took $\sim 2$mins.
@@
Figure 3: Video denoising results. Factor graphs with learnable components outperform a hand-specified pairwise smoother. Continual learning of parameters over the video further improves the PSNR over per-frame learning. Shading is ±$\pm 1$ standard error (SE) of $10$ seeds.
We see a clear benefit of depth, with significantly higher PSNR scores for the five layer model (orange/green dotted) over the single layer (orange/green dash). Despite the single layer model having only four filters, we find that it sometimes learns to reconstruct noise when learning continually over the video. To counteract this, we set the prior for each frame to be an interpolation between the previous posterior and the original prior (plotted; see App. D). In contrast, the deeper model does not require additional regularisation in this context. These results imply i) deep factor graphs have better inductive biases, capturing higher-level patterns in the data; ii) GBP Learning is effective in finding aligned multi-layer representations with only local message updates.
@@ 5.3 Image Classification
Next, we assess the our method in a supervised learning context, evaluating it on MNIST444yann.lecun.com/exdb/mnist/. MNIST has Creative Commons Attribution-Share Alike 3.0 license. The goal here is to gauge both sample efficiency and the ability to learn continually with only one pass through the training set.
@@
We train a convolutional factor graph with architecture similar to Fig. 1: a convolutional layer (with energy as per (11)), followed by a max-pool (14), dense layer (13) and (at train time) a class observation factor (16) on the output logit variables. Training is minibatched via the Bayesian filtering approach of Section 3.4, so the model only sees each datapoint once after which it can be discarded. To assess sample efficiency, we train on subsets of varying sizes, as well as the full training set. We compare against two baselines:
  1. 1.
    a linear classifier baseline which predicts logits via a dense projection of the image vector, trained with Adam over multiple epochs, and
  2. 2.
    the CNN equivalent of our factor graph, trained with Adam (kingma2014adam) for a single epoch, both i) with a FIFO replay buffer (varying sizes) and ii) without replay.
The first baseline allows us to verify our method can do nonlinear learning on images. Further, varying the size of the replay buffer in the second baseline gives an indication of how much information our model can fuse via filtering. All models were tuned on validation sets generated by randomly subsampling $15\%$ of the training set. Further details of the models and hyperparameter selection are included in App. F.
In the low data regime, GBP Learning comprehensively outperforms all baselines (Fig. 4), likely due to regularising priors and marginalising over uncertain activations. With the full training set, GBP Learning achieves a test accuracy of $98.16\pm 0.03\%$, significantly higher than the linear classifier. Further, we outperform most CNN + Adam variants except those with large replay buffers ($3600$ examples/$6\%$ of training data, $6000$ examples/$10\%$ training data), to which we perform similarly. We are thus encouraged that our model can learn complex relationships incrementally, consolidating new examples into its parameter posterior online.
Figure 4: Single epoch MNIST results. GBP Learning outperforms other methods in the small data regime, and performs similarly to a CNN with a replay buffer of $6\times 10^{3}$ examples on the full training set. Error bars cover ±$\pm 1$SE over $5$ seeds.
@@
Full dataset training with GBP Learning takes $\sim 3$ hours on NVIDIA RTX3090 GPU. While this is considerably slower than the few minutes to train the CNNs on CPU, we note that the software to train such models has benefited from decades of optimisation, as have well-established GPU hardware platforms. In contrast, our current GBP Learning implementation runs on a hardware and software stack optimised for DL. We believe orders of magnitude efficiency could be gained by running GBP Learning on a more tailored setup including processors with on-chip memory. See Section 6 for discussion.
5.3.1 Asynchronous Training
To demonstrate the robustness to different update schedules, we also trained in a layer-wise asynchronous manner. At each step, a layer is sampled uniformly, and their messages updated. Such an approach yields a final accuracy of $98.11\pm 0.04\%$ (±$\pm 1$SE over $5$ seeds), close to $98.16\pm 0.03\%$ achieved by the same model trained with synchronous forward/backward sweeps (details and plot of convergence in App. F.4). This result implies that GBP Learning can work well when the model is distributed over multiple processors, without the need for global synchronisation.
5.3.2 Comparison with lucibello2022deep(lucibello2022deep)
To understand how our method compares to previous work, we evaluate on the same image classification benchmarks presented in lucibello2022deep(lucibello2022deep): MNIST, FashionMNIST and CIFAR10. lucibello2022deep(lucibello2022deep) only support dense models where our method works with general architectures. We thus compare the convolutional model described in Section 5.3 with the dense factor graph results from lucibello2022deep(lucibello2022deep). For GBP Learning, we retain the same architecture and hyperparameters used for MNIST — we do not conduct any task-specific tuning.
The results are given in Table 1. The approaches perform similarly for FashionMNIST, but our method achieves higher accuracy for both MNIST and CIFAR10. In the latter case, we outperform lucibello2022deep(lucibello2022deep) by more than $10\%$. Moreover, we obtain such performance with incremental training, passing over the training set once only, where lucibello2022deep(lucibello2022deep) train for $100$ epochs.
[…]
Dataset lucibello2022deep(lucibello2022deep) Ours
MNIST $97.40\pm 0.04$ $\mathbf{98.16\pm 0.03}$
FashionMNIST $88.2\pm 0.1$ $88.2\pm 0.1$
CIFAR10 $41.3\pm 0.1$ $\mathbf{53.1\pm 0.3}$
6 Conclusion
@@
We have introduced GBP Learning, a method for learning in Gaussian factor graphs. Parameters are included as variables in the graph and learnt using the same BP inference procedure used to estimate all other latent variables. Inter-layer factors encourage representations to be locally consistent, and multiple layers can be stacked in order to learn richer abstractions. Experimentally, we have shown a shallow, learnable model improves over a hand-crafted method on a video denoising task, with further performance gains coming from stacking multiple layers. We have also trained convolutional factor graphs for image classification with GBP Learning. On MNIST, we demonstrate encouraging sample efficiency, and reach comparable single epoch performance to a CNN with a replay buffer of $6\times 10^{3}$ examples. GBP Learning outperforms lucibello2022deep(lucibello2022deep) by $11.8\%$ on CIFAR10 and $0.8\%$ on MNIST.
@@
While we find these initial results encouraging, we highlight scaling up GBP Learning to bigger, more complex models and datasets as an exciting future direction. Our current implementation is built on software and run on hardware heavily optimised for DL. Scaling up GBP Learning will require a bespoke hardware/software system, which can better leverage the distributed nature of GBP inference. In particular, we believe processors with memory local to the compute cores (Graphcore; cerebras) to be a promising platform for BP. Different sections of the factor graphs could be mapped to different cores. Local message updates, between factors and variables on the same core, would be cheap and could be updated at high frequency, with occasional inter-core communication to ensure alignment between different parts of the model. Low-level GBP primitives, written in e.g. Poplar555graphcore.ai/products/poplar, could be used to ensure the message updates make best use of high-bandwidth, local memory. A similar system was applied to bundle adjustment problems with GBP (Ortiz:etal:CVPR2020), and was found to be around 24×24$\times$24 × faster on IPU (Graphcore) than a state-of-the-art CPU solver.
xmain.bbl
Appendix A Energy Functions for Additional Factors
In Section 3.1, we described the energies for factors connecting convolutional and transposed convolutional layers. Here, we provide the energies for additional factors we use in our models.
For an input patch from channel $c$, $\mathbf{X}_{l-1}^{(a,b,c)}\in\mathbb{R}^{K\times K}$, centred at $(a,b)$, a connected max-pooling factor has energy:
To increase the spatial extent of a representation from layer $l$ to its input $l-1$, we use upsampling layers. These factors connect a single activation variable in $l$, $x_{l}^{(a,b,c)}$ to the patch within its receptive field in $l-1$, $\mathbf{X}_{l-1}^{(a,b,c)}\in\mathbb{R}^{K\times K}$. Their energy is
$E_{\mathrm{upsample}}^{(a,b,c)}$ $=\frac{1}{2\sigma_{l}^{2}}\sum_{i,j=-\lfloor K/2\rfloor}^{\lfloor K/2\rfloor}\left(x_{l-1}^{(a-i,b-j,c)}-x_{l}^{(a,b,c)}\right)^{2}~{}.$ (15)
Last, class supervision may then be incorporated by treating last layer activations $\mathbf{x}_{L}$ as logits and connecting them to a observation factor with energy
$E_{\mathrm{softmax}}$ $=\frac{\left\lVert\mathrm{softmax}\left(\mathbf{x}_{L}\right)-\mathbbm{1}_{y}\right\rVert_{2}^{2}}{2\sigma_{\mathrm{softmax}}^{2}}~{},$ (16)
where $\mathbbm{1}_{y}$ is a one-hot encoding of the observed class $y$.
Appendix B Factor to Variable Message Update Optimisation